# Ab initio uncertainty quantification in scattering analysis of microscopy

### v2

@@ IV.4 Model selection and diagnostics
Given a few plausible models of ISF, how do we know which one shall be used? The conventional way in DDM is to evaluate the “fit” of image structure function. However, as each transformed datum in image structure function exhibits different variances and are correlated, it may be hard to select a metric to evaluate the fit. Instead, one may focus on the “fit” in the original Cartesian space. Statistical information criteria, such as Akaike information criterion (AIC) [69], may be computed to evaluate the fit in original space with the maximum likelihood estiamtor: $\mbox{AIC}=2p-\mbox{ln}(\mathcal{L}(\bm{\theta}_{est},\mathbf{A}_{1:J,est},\bar{B}_{est}))$, where $\mathcal{L}(\bm{\theta}_{est},\mathbf{A}_{1:J,est},\bar{B}_{est})$ denotes the maximum likelihood value and $p$ is the number of parameters to be estimated. AIC quantifies the predictive error, and hence one selects a model with a small AIC. In practice, models with a larger set of parameters are often selected by AIC, when the number of observations is large. One can compensate by using Bayesian information criterion (BIC) [70], which also penalizes for the number of observations for model selection.
@@
where for isotropic processes, the $j^{\prime}\in S_{j}$ column of $\mathbf{z}_{pred}(t_{n^{*}})$ is $z_{pred,j^{\prime}}(t_{n^{*}})=A_{est,j}\mathbf{r}^{T}_{est}(t_{n^{*}})(A_{est,j}\mathbf{R}_{est,j}+\bar{B}_{est,j})^{-1}(\hat{\mathbf{y}}_{re,j^{\prime}}(\mathbf{t}_{1:(n^{*}-1)})+i\hat{\mathbf{y}}_{im,j^{\prime}}(\mathbf{t}_{1:(n^{*}-1)}))$ with $\mathbf{r}_{est}(t_{n^{*}})=(f_{\bm{\theta}_{est}}(q_{j},t_{n^{*}}-t_{1}),...,f_{\bm{\theta}_{est}}(q_{j},t_{n^{*}}-t_{n^{*}-1}))^{T}$ is an $n^{*}-1$ vector of ISF, and $\mathbf{R}_{est,j}$ being a $({n^{*}}-1)\times({n^{*}}-1)$ matrix, the $(k,k^{\prime})$th entry of which is $f_{\bm{\theta}_{est}}(q_{j},t_{k}-t_{k^{\prime}})$. On the other hand, the $j$th column of the predictive mean for anisotropic processes follows $z_{pred,j}(t)=A_{est,j}\mathbf{r}^{T}_{est}(t_{n^{*}})(A_{est,j}\mathbf{R}_{est,j}+\bar{B}_{est,j})^{-1}(\hat{\mathbf{y}}_{re,j}(\mathbf{t}_{1:(n^{*}-1)})+i\hat{\mathbf{y}}_{im,j}(\mathbf{t}_{1:(n^{*}-1)}))$, as the ISF is different for each wavevector in general. The generalized Schur algorithm is used for accelerating the computation of predictive mean in Eq. (13) to avoid the direct inversion of the covariance matrices. Then we select the model that minimizes the predictive loss, such as the average root mean squared error. When predictive errors are similar, a model with a smaller number of parameters is preferred.
@@ V.1 Isotropic processes
Figure 2: Estimating mean-squared displacements versus lag time for simulated videos containing the trajectories of $M$ = 50 simulated particles by AIUQ with reduced wavevectors (blue crosses), AIUQ all q (cyan crosses), $D(q,\Delta t)$ fit 1 (purple diamonds) and $D(q,\Delta t)$ fit 2 (yellow triangle). The analysis is for (a) Slow Brownian Motion, (b) Fast Brownian Motion, (c) Fractional Brownian Motion with subdiffusive dynamics, (d) Fractional Brownian Motion with superdiffusive motion, (e) Ornstein–Uhlenbeck process and (f) a mixture of the OU process and FBM. The solid lines denote the truth. The shaded area denotes the 95% interval for AIUQ with reduced wavevectors. The truth overlaps with both AIUQ approaches.
@@
We first compare the estimation accuracy of ISF for isotropic processes through simulations. We simulated videos from six different processes: Brownian motion (BM) with two diffusion coefficients, fractional Brownian motion (FBM) with two power parameters corresponding to subdiffusive and superdiffusive processes, respectively, Ornstein–Uhlenbeck (OU) processes, and a mixture of the OU process and FBM (OU$+$FBM). The ISFs of these processes are provided in Table 3 in Appendix A. For each process, we test the performance of each approach using both a small-sized video with $100\times 100$ pixels and $100$ time frames, and a regular-sized video with $500\times 500$ pixels and $500$ time frames. The algorithms are applicable to videos with different spatial and temporal lengths. The smaller videos contain fewer frames but retain the same timestep, particle size, and essentially the same dynamics.
@@
We compare the AIUQ approaches with a reduced number of wavevectors, and with all wavevectors, two DDM approaches, and MPT. We include an AIUQ approach uses all $J$ wavevectors. The computation can be improved by reducing the first $J_{0}$ sets (or rings) of wavevectors such that $\sum^{J_{0}}_{j=1}A_{j}/\sum^{J}_{j=1}A_{j}\geq 1-\varepsilon$ and $J_{0}/J>\beta$ with $\varepsilon=0.001$ and $\beta=0.5$, which leads to at least first $50\%$ of rings of wavevectors being selected. The same choice is applied to all simulations discussed in Sec. V.1-V.3. We found that using a less conservative threshold $\varepsilon=0.01$ and $\beta=0$ leads to around the first $25\%$ of the wavevectors being selected, nonetheless, doing so yields the same estimation for physical parameters in the intermediate scattering function. However, the estimation of the noise parameter may not be as good for small videos. Here, we reduce the number of wavevectors as an approximation to AIUQ with all wavevectors only to accelerate computation, which is distinct to other approaches that require specifying the wavevector range for estimation. The two routinely used DDM approaches have been described in Sec. III.3. Finally, in MPT, we input the known particle radius which enables the algorithm to effectively choose the optimal band-pass filter, whereas other parameters such as search radius and brightness of the center pixel need to be tuned depending on the specific case.
@@
We report estimation of the noise parameter and the truth in Table S3 in SI. DDM fit 1 produces a reasonably good estimate of the noise parameter, yet the performance in estimating the physical parameters in ISFs is not satisfying. The estimation of DDM methods can be improved by selecting a wavevector range in a case-by-case manner, and yet there is no principled way to select the wavevector range. We found that even in the diffusive case, identifying an appropriate q-range is non-trivial, as shown in Fig 1 in Sec. III.3 and Fig. 9 in Appendix B. Avoiding selecting the range of wavevector can automate the scattering analysis of video microscopy, a key feature for such analysis to be integrated into any high-throughput experiments. This has been achieved by the AIUQ approach, where utilizing information in the entire q range leads to good performance.
The true MSD, estimated MSD by MPT, two DDM and AIUQ approaches of six simulated processes with a regular video size are plotted in Fig. 2. Estimation from the two AIUQ approaches and the true MSD curves overlap for all cases, indicating precise estimation of MSD by the AIUQ approaches. Furthermore, the $95\%$ confidence interval by AIUQ (shown as the blue shaded area) is relatively short, in particular for the two BMs, but they cover most of the underlying truth for MSD. Lastly, MPT is reasonably accurate for most scenarios with case-specific tuning parameters. However, it has a noticeable discrepancy at small $\Delta t$ for BM with slower dynamics (Fig. 2(a)), and at large $\Delta t$ for almost all processes, whereas AIUQ approaches are robust at small $\Delta t$. Furthermore, we record the root mean squared error (RMSE) between the true MSD and estimated MSD by different approaches in Table S1 in SI, and the AIUQ approaches have the smallest estimation error than other approaches in all scenarios, and in particular, the RMSE by two AIUQ approaches are both 5-10 times smaller than the ones by MPT. This means that if the model is properly selected, the AIUQ approach can be more accurate than MPT in terms of estimating the MSD, and they do not require tuning parameters. Here, we must assume a certain form of the ISF, similar to most studies in DDM. In MPT, it is not necessary to prescribe such a form, though eventually one must connect the estimations derived from MPT to a physical model. Therefore, extending the AIUQ approach to encompass nonparametric estimation of ISFs [26, 44] represent a promising avenue for further exploration.
@@ V.2 Model selection
Figure 3: Predictive accuracy as demonstrated by the Root Mean Square Error (RMSE) for (a) Brownian Motion ($\sigma^{2}_{BM}=0.02$) (b) Fractional Brownian Motion ($\sigma^{2}_{FBM}=8$, $\alpha=0.6$) and (c) Ornstein–Uhlenbeck process ($\sigma^{2}{OU}=64$, $\rho=0.5$) trained on the first 300 lag times. The horizontal lines are the average of the RMSEs of each method.
@@
The predictive RMSEs of three models on held-out time frames are plotted as histograms for simulated BM, FBM, and OU processes in Fig. 3 (a)-(c), where each histogram of RMSE is averaged over every 10 time frames. The overall AvgRMSEs are shown as the horizontal lines. First, when the underlying process is BM (Fig. 3(a)), the three models have almost the same RMSE at all different frames. This is because BM is a special case of FBM with $\alpha=1$. The estimated power parameter yields $\alpha_{est}\approx 0.97$, meaning the FBM model approximately reduces to the BM model in this case. For the OU process, using the binomial approximation, one has $\sigma^{2}_{OU}(1-\rho^{\Delta t})\approx\sigma^{2}_{OU}(1-\rho)\Delta t$, when $1-\rho$ is close to zero. Thus, OU can also approximate BM relatively well when $\sigma^{2}_{OU}(1-\rho)\approx\sigma^{2}_{BM}$. Indeed, we found $\sigma^{2}_{est,OU}(1-\rho_{est})\approx 0.0196$, a value close to the true sampling model from BM with the parameter $\sigma^{2}_{BM}=0.020$. When the predictive error of these three models is similar, the preferred model is BM due to its simplicity, as it contains fewer fitting parameters in its ISF. Second, when the true process is FBM (Fig 3 (b)), the fitted FBM model consistently yields a smaller predictive RMSE compared to the other two models across most held-out time frames. The AvgRMSE indicated by the pink solid line for FBM remains identifiably lower than the other misspecified models. Third, when the true process is OU (Fig 3 (c)), the AvgRMSE of fitting an OU, plotted as the red solid line, is the smallest among the three models. Thus, we correctly select the true sampling models for all three cases using the predictive error.
Performing model selection with one-step-ahead predictive error requires sequentially predicting intensities on $n-n_{0}$ time frames, which is time-consuming. Instead, we can compute AIC based on the maximum marginal likelihood value. A model with a smaller AIC is preferred as it indicates a smaller prediction error. In Table 2, we show AIC by different models in each column. We notice that the AIC of the three models is almost identical when the underlying dynamics follow a BM, as FBM and OU approximate BM with the estimated parameter. The BM is preferred when AIC is similar, as it has a smaller number of parameters. Similar to the findings in Fig. 3, we found that the correct model has the smallest AIC, corresponding to the best fit, for FBM and OU. Hence, all true models are correctly selected by the model selection criterion.
@@ V.3 Anisotropic processes
Figure 4: Parameter estimation for anisotropic processes as shown in bar graphs. The top row shows small videos (100 $\times$ 100 $\times$ 100), and the bottom row shows regular videos (500 $\times$ 500 $\times$ 500). (a)(d) Shows the analysis of particles diffusing with two distinct diffusion coefficients along x and y, whereas (b)(c)(e)(f) show fractional Brownian motion with distinct coefficients and exponents.
Figure 5: Analyzing simulated studies with anisotropic dynamics. The mean-squared displacements versus lag time from the trajectories of $M$ = 50 simulated particles are analyzed using AIUQ (blue crosses), AIUQ all q (cyan circles and triangles), MPT (pink solid lines). The analysis is for (a) Brownian Motion, (b)(c) Fractional Brownian motion. The shaded area represents the 95% interval of AIUQ reduced q. The truth overlaps with both AIUQ approaches.
@@
We plot estimated MSD from the AIUQ and MPT in Fig. 5. In panel (a), our approach correctly identifies that the diffusion coefficient along x is 4 times as large as that along y. Panel (b) shows a subdiffusive case, where the motion along y has both a higher exponent and prefactor, causing the difference of MSD along x and y to grow with lag time. Panel (c) shows another subdiffusive scenario along both axes, where the motion along $y$ has a smaller prefactor yet a larger exponent, causing the MSD along y to grow faster and eventually approach that along x. The estimated MSDs are shown for two AIUQ approaches and MPT. Across all cases, AIUQ accurately replicates MSDs of the simulated dynamics, even when the MSDs in both directions approach one another in the third scenario. Routinely used MPT algorithms [13] typically link particles between two time frames within a pre-specified radius, which may not be optimal for anisotropic processes. Indeed, we found that identifying a suitable set of tuning parameters for MPT is harder for anisotropic processes. In comparison, one does not need to tune parameters in AIUQ approaches, and the estimated MSD from AIUQ all overlap with the truth shown in Fig. 5. Furthermore, for Fig. 5(a)-(b), MPT has a noticeable discrepancy in estimating the MSD at small $\Delta t$ due to the difficulty in separating small signal from noise, whereas AIUQ approaches are accurate. For all processes, the RMSE of estimating MSD by the AIUQ approaches is a few times smaller than the ones by MPT shown in Table S2 in SI.
DDM approaches were used for analyzing the anisotropic processes [57, 58, 7], but existing approaches only use a fraction of the transformed data along the x and y coordinates, selected by researchers, and packages in DDM do not support estimating anisotropic processes. Estimation of anisotropic processes has been implemented in our AIUQ packages [32, 33]. We utilize information from all wavevectors, yielding better accuracy in estimation and enabling appropriate uncertainty assessment without the need for selecting and additional weighing of the information from different wavevectors.
We expect that this algorithm will have broader applicability to analyzing a plethora of biological scenarios involving collective cell motion under various conditions such as chemotaxis [71], durotaxis [72], and haptotaxis [73], especially for dense settings, typically seen in a confluent cell monolayer. In Sec. VI.4, for instance, we will introduce the feasibility of analyzing the anisotropic motion for probes embedded in a lyotropic liquid crystal.

### v4

@@ IV.4 Model selection and diagnostics
Given a few plausible models of ISF, how do we know which one shall be used? The conventional way in DDM is to evaluate the “fit” of the image structure function. However, as each transformed datum in the image structure function exhibits different variances and is correlated, it may be hard to select a metric to evaluate the fit. Instead, one may focus on the “fit” in the original Cartesian space. Statistical information criteria, such as Akaike information criterion (AIC) [71], may be computed to evaluate the fit in original space with the maximum likelihood estimator: $\mbox{AIC}=2p-\mbox{log}(\mathcal{L}(\bm{\theta}_{est},\mathbf{A}_{1:J,est},\bar{B}_{est}))$, where $\mathcal{L}(\bm{\theta}_{est},\mathbf{A}_{1:J,est},\bar{B}_{est})$ denotes the maximum likelihood value and $p$ is the number of parameters to be estimated. AIC quantifies the predictive error, and hence one selects a model with a small AIC. In practice, models with a larger set of parameters are often selected by AIC, when the number of observations is large. One can compensate by using the Bayesian information criterion (BIC) [72], which also penalizes for the number of observations for model selection.
@@
where for isotropic processes, the $\mathbf{j}^{\prime}\in\mathcal{S}_{j}$ column of $\mathbf{z}_{pred}(t_{n^{*}})$ is
$z_{pred,\mathbf{j}^{\prime}}(t_{n^{*}})$ $=\mathbf{r}^{T}(t_{n^{*}})(\mathbf{R}_{j}+\eta_{est}\mathbf{I}_{n^{*}-1})^{-1}\hat{\mathbf{y}}_{\mathbf{j}^{\prime}}(\mathbf{t}_{1:(n^{*}-1)}),$
with $\mathbf{r}(t_{n^{*}})=(f_{\bm{\theta}_{est}}(q_{j},t_{n^{*}}-t_{1}),...,f_{\bm{\theta}_{est}}(q_{j},t_{n^{*}}-t_{n^{*}-1}))^{T}$ is an $n^{*}-1$ vector of ISF, and $\mathbf{R}_{j}$ is a $({n^{*}}-1)\times({n^{*}}-1)$ matrix with the $(k,k^{\prime})$th entry of being $f_{\bm{\theta}_{est}}(q_{j},t_{k}-t_{k^{\prime}})$, and $\eta_{est}=\bar{B}_{est,j}A^{-1}_{est,j}$. The generalized Schur algorithm is used for accelerating the computation of predictive mean in Eq. (13) to avoid the direct inversion of the covariance matrices. Then we select the model that minimizes the predictive loss, such as the average root mean squared error. When predictive errors are similar, a model with a smaller number of parameters is preferred.
@@ V.1 Isotropic processes
Figure 2: Estimated mean-squared displacements versus lag time plotted in log-log space for simulated videos containing the trajectories of $M$ = 50 simulated particles by AIUQ with reduced wave vectors (blue crosses), AIUQ all q (cyan crosses), $D(q,\Delta t)$ fit 1 (purple diamonds) and $D(q,\Delta t)$ fit 2 (yellow triangle). The analysis is for (a) Slow Brownian Motion, (b) Fast Brownian Motion, (c) Fractional Brownian Motion with subdiffusive dynamics, (d) Fractional Brownian Motion with superdiffusive motion, (e) Ornstein–Uhlenbeck process and (f) a mixture of the OU process and FBM. The solid lines denote the truth. The shaded area denotes the 95% interval for AIUQ with reduced wave vectors. The truth overlaps with both AIUQ approaches.
@@
We first compare the estimation accuracy of ISF for isotropic processes through simulations. We simulated videos from six different processes: Brownian motion (BM) with two diffusion coefficients, fractional Brownian motion (FBM) with two power parameters corresponding to subdiffusive and superdiffusive processes, respectively, Ornstein–Uhlenbeck (OU) processes, and a mixture of the OU process and FBM (OU$+$FBM). The ISFs of these processes are provided in Table 3 in Appendix A. For each process, we test the performance of each approach using both a small-sized video with $100\times 100$ pixels and $100$ time frames, and a regular-sized video with $500\times 500$ pixels and $500$ time frames. The algorithms are applicable to videos with different spatial and temporal lengths. The smaller videos contain fewer frames but retain the same dynamics.
@@
We include the AIUQ approaches with all wave vectors, reduced wave vectors, two DDM approaches with all wave vectors, and MPT. The first AIUQ approach uses all wave vectors, which is the default setting in our package. The second AIUQ approach accelerates the computation by reducing the first $J_{0}$ sets (or rings) of wave vectors such that $\sum^{J_{0}}_{j=1}A_{j}/\sum^{J}_{j=1}A_{j}\geq 1-\varepsilon$ and $J_{0}/J>\beta$ with $\varepsilon=0.001$ and $\beta=0.5$, which leads to at least first $50\%$ of rings of wave vectors being selected. The same choice is applied to all simulations discussed in Sec. V.1-V.3. We found that using a less conservative threshold $\varepsilon=0.01$ and $\beta=0$ leads to around the first $25\%$ of the wave vectors being selected, nonetheless, doing so yields the same estimation for physical parameters in the intermediate scattering function. However, the estimation of the noise parameter may not be as good for small videos. Here, we reduce the number of wave vectors as an approximation to AIUQ with all wave vectors only to accelerate computation, which is distinct to other approaches that require specifying the wave vector range for estimation. The default setting of our packages is to use all wave vector ranges. The two routinely used DDM approaches have been described in Sec. III.4. Finally, in MPT, we input the known particle radius which enables the algorithm to effectively choose the optimal band-pass filter, whereas other parameters such as search radius and brightness of the center pixel need to be tuned depending on the specific case.
@@
We report estimation of the noise parameter and the truth in Table S3 in the Supplemental Material. DDM fit 1 produces a reasonably good estimate of the noise parameter, yet the performance in estimating the physical parameters in ISFs is not satisfying. The estimation of DDM methods can be improved by selecting a wave vector range in a case-by-case manner, and a principled way to select the wave vector range is yet to be found. We found that even in the diffusive case, identifying an appropriate q-range is non-trivial, as shown in Fig 1 in Sec. III.3 and Fig. 9 in Appendix B. Avoiding selecting the range of wave vector can automate the scattering analysis of video microscopy, a key feature for such analysis to be integrated into any high-throughput experiment. This has been achieved by the AIUQ approach, where utilizing information in the entire q range leads to good performance.
The true MSD, estimated MSD by MPT, two DDM and AIUQ approaches of six simulated processes with a regular video size are plotted in Fig. 2. Estimation from the two AIUQ approaches and the true MSD curves overlap for all cases, indicating precise estimation of MSD by the AIUQ approaches. Furthermore, the $95\%$ confidence interval by AIUQ (shown as the blue shaded area) is relatively short, in particular for the two BMs, but they cover most of the underlying truth for MSD. Lastly, MPT is reasonably accurate for most scenarios with case-specific tuning parameters. However, it has a noticeable discrepancy at small $\Delta t$ for BM with slower dynamics (Fig. 2(a)), and at large $\Delta t$ for almost all processes, whereas AIUQ approaches are robust at small $\Delta t$. Furthermore, we record the root mean squared error (RMSE) between the true MSD and estimated MSD by different approaches in Table S1 in the Supplemental Material. and the AIUQ approaches have the smallest estimation error than other approaches in all scenarios, and in particular, the RMSE by two AIUQ approaches are both 5-10 times smaller than the ones by MPT. This means that if the model is properly selected, the AIUQ approach can be more accurate than MPT in terms of estimating the MSD, and they do not require tuning parameters. For all methods, estimates for $A(q)$ are presented in Fig. S1. Here, we must assume a certain form of the ISF, similar to most studies in DDM. In MPT, it is not necessary to prescribe such a form, though eventually one must connect the estimations derived from MPT to a physical model. Therefore, extending the AIUQ approach to encompass nonparametric estimation of ISFs [27, 45] represents a promising avenue for further exploration.
@@ V.2 Model selection
Figure 3: Predictive accuracy as demonstrated by the Root Mean Square Error (RMSE) for (a) Brownian Motion ($\sigma^{2}_{BM}=0.02$) (b) Fractional Brownian Motion ($\sigma^{2}_{FBM}=8$, $\alpha=0.6$), and (c) Ornstein–Uhlenbeck process ($\sigma^{2}_{OU}=64$, $\rho=0.5$) trained on the first 350 lag times. The horizontal lines are the average of the RMSEs of each method. The RMSEs are averaged every 10 lag times to improve the readability of the graph.
@@
The predictive RMSEs of three models on held-out time frames are plotted as histograms for simulated BM, FBM, and OU processes in Fig. 3 (a)-(c), where each histogram of RMSE is averaged every 10 lag times. The overall AvgRMSEs are shown as the horizontal lines. First, when the underlying process is BM (Fig. 3(a)), the three models have almost the same RMSE at all different frames. This is because BM is a special case of FBM with $\alpha=1$. The estimated power parameter yields $\alpha_{est}\approx 0.97$, meaning the FBM model approximately reduces to the BM model in this case. For the OU process, using the binomial approximation, one has $\sigma^{2}_{OU}(1-\rho^{\Delta t})\approx\sigma^{2}_{OU}(1-\rho)\Delta t$, when $1-\rho$ is close to zero. Thus, OU can also approximate BM relatively well when $\sigma^{2}_{OU}(1-\rho)\approx\sigma^{2}_{BM}$. Indeed, we found $\sigma^{2}_{est,OU}(1-\rho_{est})\approx 0.0196$, a value close to the true sampling model from BM with the parameter $\sigma^{2}_{BM}=0.020$. When the predictive error of these three models is similar, the preferred model is BM due to its simplicity, as it contains fewer fitting parameters in its ISF. Second, when the true process is FBM (Fig 3 (b)), the fitted FBM model consistently yields a smaller predictive RMSE compared to the other two models across most held-out time frames. The AvgRMSE indicated by the pink solid line for FBM remains identifiably lower than the other misspecified models. Third, when the true process is OU (Fig 3 (c)), the AvgRMSE of fitting an OU, plotted as the red solid line, is the smallest among the three models. Thus, we correctly select the true sampling models for all three cases using the predictive error.
Performing model selection with one-step-ahead predictive error requires sequentially predicting intensities on $n-n_{0}$ time frames, which is time-consuming. Instead, we can compute AIC based on the maximum marginal likelihood value. A model with a smaller AIC is preferred as it indicates a smaller prediction error. In Table 2, we show AIC by different models in each column. We notice that the AIC of the three models is almost identical when the underlying dynamics follow a BM, as FBM and OU approximate BM with the estimated parameter. The BM is preferred when AIC is similar, as it has a smaller number of parameters. Similar to the findings in Fig. 3, we found that the correct model has the smallest AIC, corresponding to the best fit, for FBM and OU. Hence, all true models are correctly selected by the model selection criterion. Automated model selection is not well-studied in scattering analysis of microscopy videos. Therefore, improving the robustness of the methods in complex experimental conditions is of great interest in practice.
@@ V.3 Anisotropic processes
Figure 4: Parameter estimation for anisotropic processes, where the truth is plotted as bars.The top row shows small videos (100 $\times$ 100 $\times$ 100), and the bottom row shows regular videos (500 $\times$ 500 $\times$ 500). (a)(d) show the analysis of particles diffusing with two distinct diffusion coefficients along x and y, whereas (b)(c)(e)(f) show fractional Brownian motion with distinct coefficients and exponents.
Figure 5: Analyzing simulated studies with anisotropic dynamics. The estimated mean-squared displacements versus lag time are plotted in log-log space based on simulation with $M$ = 50 simulated particles, by AIUQ (blue crosses), AIUQ all q (cyan circles and triangles), MPT (pink solid lines). The analysis is for (a) Brownian Motion, (b)(c) Fractional Brownian motion. The shaded area represents the 95% interval of AIUQ reduced q. The truth overlaps with both AIUQ approaches.
@@
We plot estimated MSD from the AIUQ and MPT in Fig. 5. In panel (a), our approach correctly identifies that the diffusion coefficient along x is 4 times as large as that along y. Panel (b) shows a subdiffusive case, where the motion along y has both a higher exponent and prefactor, causing the difference of MSD along x and y to grow with lag time. Panel (c) shows another subdiffusive scenario along both axes, where the motion along $y$ has a smaller prefactor yet a larger exponent, causing the MSD along y to grow faster and eventually approach that along x. The estimated MSDs are shown for two AIUQ approaches and MPT. Across all cases, AIUQ accurately replicates MSDs of the simulated dynamics, even when the MSDs in both directions approach one another in the third scenario. Routinely used MPT algorithms [13] typically link particles between two time frames within a pre-specified radius, which may not be optimal for anisotropic processes. Indeed, we found that identifying a suitable set of tuning parameters for MPT is harder for anisotropic processes. In comparison, one does not need to tune parameters in AIUQ approaches, and the estimated MSD from AIUQ all overlap with the truth shown in Fig. 5. Furthermore, for Fig. 5(a)-(b), MPT has a noticeable discrepancy in estimating the MSD at small $\Delta t$ due to the difficulty in separating small signal from noise, whereas AIUQ approaches are accurate. For all processes, the RMSE of estimating MSD by the AIUQ approaches is a few times smaller than the ones by MPT shown in Table S2 in the Supplemental Material. .
DDM approaches were used for analyzing the anisotropic processes [59, 60, 7], but existing approaches only use a fraction of the transformed data along the x and y coordinates, selected by researchers, and existing DDM packages do not support estimating anisotropic processes. Estimation of anisotropic processes has been implemented in our AIUQ packages [33, 34]. We utilize information from all wave vectors, yielding better accuracy in estimation and enabling appropriate uncertainty assessment without the need for selecting and additional weighing of the information from different wave vectors.
We expect that this algorithm will have broad applicability to analyzing a plethora of biological scenarios involving collective cell motion under various conditions such as chemotaxis [73], durotaxis [74], and haptotaxis [75], especially for dense settings, typically seen in a confluent cell monolayer. In Sec. VI.4, for instance, we will introduce the feasibility of analyzing the anisotropic motion for probes embedded in a lyotropic liquid crystal.